# Recovering 21cm monopole signals without smoothness

### v3

@@ 2 Separating foregrounds
where we have split the signal into monopole signal of interest $T_{\rm s}(\nu)$ and the foreground temperature $T_{\rm fg}(\hat{\mathbf{x}},\nu)$. In this treatment, we ignore spatial fluctuations in the cosmological signal and focus solely on the monopole signal. The fluctuations in the foreground are typically several orders of magnitude brighter than the signal. We further split them into the mean foreground monopole $\bar{T}_{\rm fg}(\nu)$ and variations around this mean $\Delta T_{\rm fg}(\hat{\mathbf{x}},\nu)$. By definition, then, these fluctuations must average to zero $\Delta T_{\rm fg}(\hat{\mathbf{x}},\nu)\right>_{\hat{\mathbf{x}}}=0$, where the average is defined as usual over a sufficiently large observational patch.
@@ 2.1 Smoothness separation
To further illustrate this point on a concrete example, we consider the following toy model: suppose there are two patches of the sky, each emitting a pure power-law emission with spectral indices -2.54 and -2.56, respectively, and contributing equally to the total emission. We fit the total mean foreground with a maximally smooth function (MSF) using the maxsmooth package (see Bevins et al. (2021)) with the basis chosen as polynomials, log-polynomials, and loglog-polynomials as the basis of maximum order 15 in the expansion for the fit. As shown in Figure 2, the best-fits still show residuals of various amplitudes that can be confused as a false detection. (We note that the log-polynomials with order N=15 have a markedly poor fit despite the large order and need significantly more terms that take an unreasonably long time to converge). The crucial point is that residuals have structures that vary over similar scales as the actual signal, while the amplitude is still over an order of magnitude larger than the signal of interest. Both the maxsmooth paper and the earlier implementations in Rao et al. (2015); Sathyanarayana Rao et al. (2017) recommend several diagnostics for improving and validating the fitting routines that could be implemented in this case. Still, the need for additional diagnostics demonstrates the central issue of separation based on foreground smoothness).
@@ 3 Method
For a start, we assume a perfect instrument without any calibration errors or gain fluctuations. In this case, assuming temperature data from the sky $T(\hat{\mathbf{x}})$, we can calculate the mean sky monopole data,
@@ 4.4 Rotation into PCA basis
Figure 4: This figure shows the separation of foregrounds and the Cosmic Dawn signal in different spaces for the Global Sky Model. Top Left: Separation in the original frequency space using an achromatic $2^{\circ}$ beam. The blue distribution violins and the solid line describe the foreground signal distribution and mean (on the cut sky), the green line denotes the CMB signal, and the orange solid line denotes the Dark Ages signal. Note the extremely large dynamic range on the vertical axis. Top Right: The amplitude of the same signals in the foreground PCA basis. The first six principal components absorb about 6 orders of magnitude of the signal. The brown solid line denotes the RMS of the corresponding eigenmode. Bottom Panel Signal amplitudes scaled by the RMS of the corresponding eigenmode. The basis vectors have been multiplied by $-1$ to make all the Cosmic Dawn contributions positive.
@@ 4.7 CMB Signal Model
Figure 5: Model approximations to the true 21cm signal used in this analysis. Both the Dark Ages and Cosmic Dawn models are 3-parameter models that depend on the amplitude, width and center frequency, and are scaled to approach zero at the boundaries of their domains.
@@
We build phenomenological 3-parameter models for the Dark Ages and Cosmic Dawn signals as follows. We start with the true 21 cm signal shape as predicted by the ARES software package by Mirocha (2014). We then scale this model in amplitude, position, and width:
@@
where $A_{0}=40\,$mK, $\nu_{c0}=16.4\,$Mhz and $\Delta\nu_{0}=14\,$Mhz are (defined to be) values of amplitude, trough center and width for the Dark Ages signal and $A_{0}=130\,$mK, $\nu_{c0}=67.5\,$Mhz and $\Delta\nu_{0}=20\,$Mhz are (defined to be) the values of amplitude, trough center and width for the Cosmic Dawn Signal. To avoid the method latching on to the mid-frequency inflection region at 50 MHz, we have manually modified both the models so that they gracefully approach zero at 50MHz by matching first and second derivatives at 32 MHz, 55 MHz and 85 MHz. See Figure 5 for the exact model shapes compared to the true 21cm signal.
@@ 4.9 Likelihood analysis
Our signal model has three degrees of freedom, and therefore, the likelihood analysis can be done by brute-force evaluation on a sufficiently fine 3D grid. After we evaluate likelihood on a regular grid, we use the standard python MCMC visualization package corner from Foreman-Mackey (2016) to plot and calculate marginalized constraints by assigning sample weights proportional to their likelihood. Where the likelihood is prior-dominated, we instead calculate the amplitude-only likelihoods by fixing the width and center to their true values, giving us at least the exclusion limits on the amplitude of the signal of interest.
@@ 5 Results
Figure 6: Cosmic Dawn signal constraints for achromatic beams for the ideal noiseless, noisy, and noiseless instrumental systematics cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. Gray areas represent the physical prior on the three parameters and dashed black indicate the true values of the parameters. All results correspond to the same fixed noise realization.
Figure 7: Dark Ages signal constraints for achromatic beams for the ideal noiseless, noisy, and noiseless instrumental systematics cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. Gray areas represent the physical prior on the three parameters and dashed black indicate the true values of the parameters. All results correspond to the same fixed noise realization.
Figure 8: Signal amplitude likelihood constraints for achromatic and chromatic beams, with the width and center fixed to the true values, for the ideal noiseless, noisy, and noiseless instrumental systematics cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. We plot achromatic (top pair, regular) and chromatic beams (bottom pair, hatched) for each combination. Gray areas represent the physical prior on the three parameters and dashed black lines indicate the true values of the parameters. All results correspond to the same fixed noise realization.
Figure 9: Signal amplitude likelihoods for the Cosmic Dawn signal and the Dark Ages signal for ten different noise realizations at different signal-to-noise levels. The solid lines represent the median likelihoods over all noise realizations, and the shaded regions correspond to the 1-sigma variation across the different noise realizations. The dashed vertical line at $40$ mK and 130 mK is the true value for the signal amplitude for the Dark Ages model and Cosmic Dawn model, respectively.
Figure 10: Signal amplitude likelihoods for the Cosmic Dawn signal and the Dark Ages signal for ten different noise realizations at different signal-to-noise levels. The solid lines represent the median likelihoods over all noise realizations, and the shaded regions correspond to the 1-sigma variation across the different noise realizations. The dashed vertical line at $40$ mK and 130 mK is the true value for the signal amplitude for the Dark Ages model and Cosmic Dawn model, respectively.
@@ 5.1 Results in an idealistic case
Our idealistic case is a perfectly calibrated, noiseless instrument. Results from the Figures 6, 7, and 8 can be summarized as follows. First, we find that in majority of cases we have detection of the signal with high confidence. The chromatic multi-scale case recovers the true values of parameters with percent level confidence. We also find that the results are unbiased, with the exception of $2^{\circ}+4^{\circ}+6^{\circ}$ achromatic Dark Ages case (discussed below), the results are consistent with the ground truth at two sigma level, even when sensitivity is at the percent level.
@@ 5.2 Effects of adding noise
We now consider the effect of adding noise to the maps. We find that for the pessimistic noise scenario (see 1), with an SNR of 108 for the Dark Ages signal and 105 for the Cosmic Dawn signal, it is possible to constrain the upper limit on the amplitude of the signal of interest with some confidence, especially in the achromatic case (see Figures 8, 9, 10). For chromatic beams, the amplitude likelihood constraints for Dark Ages multi-scale fits are biased towards higher values and inconsistent with the true amplitude. We find that independent noise realizations can lead to different values of this bias. However, the sizes of the errors remain consistent (see Appendix B for details on systematic and noise-realization biases). The three-parameter likelihoods are prior-dominated and do not provide any useful constraints at this SNR level.
@@ 5.4 Detecting CMB
Figure 11: CMB Signal constraints for achromatic(filled box plots) and chromatic (hatched box plots) beams for the ideal noiseless and noisy cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. Dashed black indicates the true values of the parameters. All results correspond to the same fixed noise realization.

### v5

@@ 2 Separating foregrounds
where we have split the signal into monopole signal of interest $T_{\rm s}(\nu)$ and the foreground temperature $T_{\rm fg}(\hat{\mathbf{x}},\nu)$.
In this treatment, we ignore spatial fluctuations in the cosmological signal and focus solely on the monopole signal. The fluctuations in the foreground are typically several orders of magnitude brighter than the signal. We further split them into the mean foreground monopole $\bar{T}_{\rm fg}(\nu)$ and variations around this mean $\Delta T_{\rm fg}(\hat{\mathbf{x}},\nu)$. By definition, then, these fluctuations must average to zero $\Delta T_{\rm fg}(\hat{\mathbf{x}},\nu)\right>_{\hat{\mathbf{x}}}=0$, where the average is defined as usual over a sufficiently large observational patch.
@@ 2.1 Smoothness separation
To further illustrate this point on a concrete example, we consider the following toy model: suppose there are two patches of the sky, each emitting a pure power-law emission with spectral indices -2.54 and -2.56, respectively, and contributing equally to the total emission. We fit the total mean foreground with a maximally smooth function (MSF) using the maxsmooth package (see Bevins et al. (2021)) with the basis chosen as polynomials, log-polynomials, and loglog-polynomials as the basis of maximum order 15 in the expansion for the fit. As shown in Figure 2, the best-fits still show residuals of various amplitudes that can be confused as a false detection. (We note that the log-polynomials with order N=15 have a markedly poor fit despite the large order and need significantly more terms that take an unreasonably long time to converge). The crucial point is that residuals have structures that vary over similar scales as the actual signal, while the amplitude is still over an order of magnitude larger than the signal of interest. Both the maxsmooth paper and the earlier implementations in Rao et al. (2015); Sathyanarayana Rao et al. (2017) recommend several diagnostics for improving and validating the fitting routines that could be implemented in this case. In particular, the use of Bayesian Information Criterion and the change in Bayesian evidence after jointly fitting the signal and foregrounds, can be used to evaluate the goodness of fit and are recommended in the literature to evaluate relative modeling systematics. Still, the need for additional diagnostics demonstrates the central issue of separation based on foreground smoothness.
@@ 3 Method
For a start, we assume a perfect instrument without any calibration errors or gain fluctuations. In this case, assuming temperature data from the beam-convolved sky $T(\hat{\mathbf{x}})$, we can calculate the mean sky monopole data,
@@ 4.4 Rotation into PCA basis
Figure 4: This figure shows the separation of foregrounds and the Cosmic Dawn signal in different spaces for the Global Sky Model. Top Left: Separation in the original frequency space using an achromatic $2^{\circ}$ beam. The blue distribution violins and the solid line describe the foreground signal distribution and mean (on the cut sky), the green line denotes the CMB signal, and the orange solid line denotes the Cosmic Dawn signal. Note the extremely large dynamic range on the vertical axis. Top Right: The amplitude of the same signals in the foreground PCA basis. The first six principal components absorb about 6 orders of magnitude of the signal. The brown solid line denotes the RMS of the corresponding eigenmode. Bottom Panel Signal amplitudes scaled by the RMS of the corresponding eigenmode. The basis vectors have been multiplied by $-1$ to make all the Cosmic Dawn contributions positive.
@@ 4.7 CMB Signal Model
Figure 5: Model approximations to the fiducial 21cm signal used in this analysis. Both the Dark Ages and Cosmic Dawn models are 3-parameter models that depend on the amplitude, width and center frequency, and are scaled to approach zero at the boundaries of their domains.
@@
We build phenomenological 3-parameter models for the Dark Ages and Cosmic Dawn signals as follows. We start with the fiducial 21 cm signal shape as predicted by the ARES software package by Mirocha (2014). We then scale this model in amplitude, position, and width:
@@
where $A_{0}=40\,$mK, $\nu_{c0}=16.4\,$Mhz and $\Delta\nu_{0}=14\,$Mhz are (defined to be) values of amplitude, trough center and width for the Dark Ages signal and $A_{0}=130\,$mK, $\nu_{c0}=67.5\,$Mhz and $\Delta\nu_{0}=20\,$Mhz are (defined to be) the values of amplitude, trough center and width for the Cosmic Dawn Signal. To avoid the method latching on to the mid-frequency inflection region at 50 MHz, we have manually modified both the models so that they gracefully approach zero at 50MHz by matching first and second derivatives at 32 MHz, 55 MHz and 85 MHz. See Figure 5 for the exact model shapes compared to the fiducial 21cm signal.
@@ 4.9 Likelihood analysis
Our signal model has three degrees of freedom, and therefore, the likelihood analysis can be done by brute-force evaluation on a sufficiently fine 3D grid ($N=70^{3}$). In both the 3D grid and the 1D amplitude likelihoods, we use flat priors for the respective parameter-spaces. For the Dark Ages 3D grids, we use the priors $A\in(0.4\,{\rm mK},400\,{\rm mK}),\nu_{\rm rms}\in(10\,{\rm MHz},30\,{\rm MHz}),\nu_{\rm min}\in(10\,{\rm MHz},30\,{\rm MHz})$, and for Cosmic Dawn we use $A\in(1.3\,{\rm mK},1300\,{\rm mK}),\nu_{\rm rms}\in(10\,{\rm MHz},40\,{\rm MHz}),\nu_{\rm min}\in(50\,{\rm MHz},90\,{\rm MHz})$ for the signal amplitude, width and center respectively. For Dark Ages, we initially tested the 3D grid limits by imposing the physical $\nu_{\rm rms},\nu_{\rm min}\in(1\,{\rm MHz},50\,{\rm MHz})$ for the both the signal width and center, which correspond to the bin width and maximum frequency which was further reduced to the limits above to improve the grid density around the regions of interest while leaving the likelihoods unchanged. For Cosmic Dawn, we similarly tested with the full $A\in(50\,{\rm MHz},100\,{\rm MHz})$ range for center, $\nu_{\rm rms}\in(10\,{\rm MHz},40\,{\rm MHz})$ range for width, and $\nu_{\rm min}\in(50\,{\rm MHz},90\,{\rm MHz})$, and reduced it to the above. For the 1D amplitude likelihoods we use the flat priors $(0.1A_{0},1000A_{0})$ where $A_{0}=40$ mK for Dark Ages and $130$ mK for Cosmic Dawn, respectively.
After we evaluate likelihood on a regular grid, we use the standard python MCMC visualization package corner from Foreman-Mackey (2016) to plot and calculate marginalized constraints by assigning sample weights proportional to their likelihood. Where the likelihood is prior-dominated, we instead calculate the amplitude-only likelihoods by fixing the width and center to their fiducial values, giving us at least the exclusion limits on the amplitude of the signal of interest.
@@ 5 Results
Figure 6: Cosmic Dawn signal constraints for achromatic beams for the ideal noiseless, noisy, and noiseless instrumental systematics cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. Gray areas represent the physical prior on the three parameters and dashed black indicate the fiducial values of the parameters. All results correspond to the same fixed noise realization.
Figure 7: Dark Ages signal constraints for achromatic beams for the ideal noiseless, noisy, and noiseless instrumental systematics cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. Gray areas represent the physical prior on the three parameters and dashed black indicate the fiducial values of the parameters. All results correspond to the same fixed noise realization.
Figure 8: Signal amplitude likelihood constraints for achromatic and chromatic beams, with the width and center fixed to the fiducial values, for the ideal noiseless, noisy, and noiseless instrumental systematics cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. We plot achromatic (top pair, regular) and chromatic beams (bottom pair, hatched) for each combination. Gray areas represent the physical prior on the three parameters and dashed black lines indicate the fiducial values of the parameters. All results correspond to the same fixed noise realization.
Figure 9: Signal amplitude likelihoods for the Cosmic Dawn signal and the Dark Ages signal for ten different noise realizations at different signal-to-noise levels for achromatic beams. The solid lines represent the median likelihoods over all noise realizations, and the shaded regions correspond to the 1-sigma variation across the different noise realizations, with an added offset to aid readability. The vertical line at $40$ mK and 130 mK is the fiducial value for the signal amplitude for the Dark Ages model and Cosmic Dawn model, respectively.
Figure 10: Signal amplitude likelihoods for the Cosmic Dawn signal and the Dark Ages signal for ten different noise realizations at different signal-to-noise levels for chromatic beams. The solid lines represent the median likelihoods over all noise realizations, and the shaded regions correspond to the 1-sigma variation across the different noise realizations, with an added offset to aid readability. The vertical line at $40$ mK and 130 mK is the fiducial value for the signal amplitude for the Dark Ages model and Cosmic Dawn model, respectively.
@@ 5.1 Results in an idealistic case
Our idealistic case is a perfectly calibrated, noiseless instrument. Results from the Figures 6, 7, and 8 can be summarized as follows. First, we find that in majority of cases we have detection of the signal with high confidence. The chromatic multi-scale case recovers the fiducial values of parameters with percent level confidence. We also find that the results are unbiased, with the exception of $2^{\circ}+4^{\circ}+6^{\circ}$ achromatic Dark Ages case (discussed below), the results are consistent with the ground truth at two sigma level, even when sensitivity is at the percent level.
@@ 5.2 Effects of adding noise
We now consider the effect of adding noise to the maps. We find that for the pessimistic noise scenario (see 1), with an SNR of 108 for the Dark Ages signal and 105 for the Cosmic Dawn signal, it is possible to constrain the upper limit on the amplitude of the signal of interest with some confidence, especially in the achromatic case (see Figures 8, 9, 10). For chromatic beams, the amplitude likelihood constraints for Dark Ages multi-scale fits are biased towards higher values and inconsistent with the fiducial amplitude. We find that independent noise realizations can lead to different values of this bias. However, the sizes of the errors remain consistent (see Appendix B for details on systematic and noise-realization biases). The three-parameter likelihoods are prior-dominated and do not provide any useful constraints at this SNR level.
@@ 5.4 Detecting CMB
Figure 11: CMB Signal constraints for achromatic(filled box plots) and chromatic (hatched box plots) beams for the ideal noiseless and noisy cases. Box plot boxes represent 1-sigma errors, median line in the center, and whiskers bound the 95% confidence intervals for each case. Dashed black indicates the fiducial values of the parameters. All results correspond to the same fixed noise realization.